1 child no children 2 children 3+ children 1 adult 6 8 2 adults 6 11 6 1.

Suppose that a food aid program provides every household in the apartment complex described in the table above with an extra $30 of food benefits per week for each adult in the household, and $10 per week per child (up to three). Find the average aid amount per household.
Answer #1

Answer)

It is given that each adult will get $30 worth food benefit.

So, in first row

We have, 5+6+8+4 = 23 house holds with one adult

So, 23*30 = $690

And in second row we have 6+11+9+6 = 32 households with two adults

So, we have 32*30 = $960

Similarly each child will get $10

In first row, we have 6 house holds with 1 children and 8 with 2 and 4 with 3

Total = 6*1 + 8*2 + 4*3 = 6+16+12 = 34

Total aid = 34*10 = 340

Similarly in second row

Total aid = 11*1+9*2+6*3 = 44

Aid = 440

Total aid = 440+340+960+690 = $2430

Total number of households = 5+6+6+11+8+9+4+6 = 55

average= 2430/55

= 44.181818181818